### v1

# The strength of a geometric simplex is a key ingredient in a polynomial-time classification of unordered point clouds by Lipschitz continuous invariants

Abstract
The basic input for many real shapes is a finite cloud of unordered points. The strongest equivalence between shapes in practice is Euclidean motion. The recent polynomial-time classification of point clouds required a Lipschitz continuous function that vanishes on degenerate simplices, while the usual volume is not Lipschitz. We define the strength of any geometric simplex and prove its continuity under perturbations with explicit bounds for Lipschitz constants.
keywords:
unordered point cloud , rigid motion , complete classification , Lipschitz continuity , geometric simplex
MSC:
51F20 , 51N20 , 51M25 , 15A15
journal: Applied Mathematics Letters
[…]
organization=School of Computer Science and Informatics, University of Liverpool,addressline=Ashton street, city=Liverpool, postcode=L69 3BX, state=, country=United Kingdom
1 The importance of Lipschitz continuity for distinguishing mirror images under noise
Many applications deal with point configurations or clouds of points obtained as edge pixels or feature points of objects across all scales from galaxies to molecules [1]. Positions of points in a Euclidean space, such as atomic centres, are always uncertain due to measurement noise or thermal vibrations. Molecular dynamics simulates trajectories of atomic clouds, while machine learning tries to predict molecular properties that depend on atomic geometry.
The robustness of algorithmic predictions means that an output only continuously changes under perturbations of given points. The classical $\varepsilon-\delta$ concept of continuity is very weak in the sense that all standard functions are continuous on domains, where they are defined. For example, $f(x)=\dfrac{1}{x}$ is continuous for any $x\neq 0$ because, for any $\varepsilon>0$, there is $\delta=\dfrac{|x|}{2}\min\{1,\varepsilon|x|\}$ such that if $|x-y|<\delta$, then $|f(x)-f(y)|=\dfrac{|x-y|}{|xy|}\leq\dfrac{\delta}{|xy|}\leq\dfrac{2\delta}{x^{2}}\leq\varepsilon$. However, for small $x>0$, the chosen delta $\delta$ is much smaller than $\varepsilon$, so $f(x)=\dfrac{1}{x}$ grows too fast close to 0. The Lipschitz continuity below is much more practical by restricting the growth of a function via a constant and an amount of perturbation.
Definition 1.1 (Lipschitz continuity).
A function $f:\mathbb{R}^{m}\to\mathbb{R}$ is Lipschitz continuous if there is a Lipschitz constant $\lambda>0$ such that $|f(x)-f(y)|\leq\lambda|x-y|$ for any $x,y\in\mathbb{R}^{n}$, where |xy||x-y| denotes the Euclidean distance.
Then $f(x)=\dfrac{1}{x}$ is not Lipschitz continuous because for any $\lambda>0$, we can set $c=\max\{1,\lambda\}$, $x=\dfrac{1}{2c}$ and $y=\dfrac{1}{c}$ such that $|f(x)-f(y)|=c\geq 1>\dfrac{\lambda}{2c}=\lambda|x-y|$. Though the Lipschitz continuity makes sense for maps between metric spaces, we consider only scalar functions on simplices in $\mathbb{R}^{n}$. A simplex can be defined as a finite set of elements whose every subset is also a simplex. We consider only geometric realizations of a simplex, still called a simplex.
Definition 1.2 (a geometric simplex $A$ on $n+1$ points in $\mathbb{R}^{n}$).
The (geometric) simplex $A$ on any $n+1$ ordered points $p_{0},\dots,p_{n}\in\mathbb{R}^{n}$ is the subset $A=\left\{\sum\limits_{i=0}^{n}t_{i}p_{i}\;\vline\;t_{i}\in[0,1],\;\sum\limits_{i=0}^{n}t_{i}=1\right\}\subset\mathbb{R}^{n}$ with ordered vertices $p_{0},\dots,p_{n}$. An orientation of $A$ is the sign of the determinant of the $n\times n$ matrix with the columns $p_{1}-p_{0},\dots,p_{n}-p_{0}$, and is denoted by $\mathrm{sign}(A)$.
In dimension $n=1$, the simplex on any points $p_{0},p_{1}\in\mathbb{R}$ is the line segment connecting $p_{0}$ and $p_{1}$. If points $p_{0},\dots,p_{n}\in\mathbb{R}^{n}$ are affinely independent, i.e. there is no (n1)(n-1)-dimensional affine subspace of $\mathbb{R}^{n}$ containing all $p_{0},\dots,p_{n}$ and hence the simplex $A$, then $A$ is $n$-dimensional. However, Definition 1.2 makes sense for any points.
The volume $\mathrm{vol}(A)$ of a simplex $A$ or an arbitrary polyhedron is often used as a shape descriptor and detects affine independence in the sense that $A$ is degenerate if and only if $\mathrm{vol}(A)=0$. However, the volume and all other distance-based descriptors do not distinguish mirror images, which have different signs of orientation.
When $A$ goes through a degenerate configuration, an orientation of $A$ can discontinuously change the sign. This discontinuity is an obstacle to recognizing simplices (or more general clouds) that are nearly mirror-symmetric.
One attempt to resolve this discontinuity is to consider the signed volume $\mathrm{sign}(A)\mathrm{vol}(A)$, because $\mathrm{vol}(A)$ vanishes only on degenerate simplices. However, $\mathrm{vol}(A)$ is not Lipschitz continuous for any $n\geq 2$, as illustrated below.
Example 1.3 (the area of a triangle is not Lipschitz continuous).
For any large $l>0$ and real $\varepsilon$ close to 0, let $T(l,\varepsilon)\subset\mathbb{R}^{2}$ be the 2D simplex (triangle) on the vertices $(0,\varepsilon)$ and $(\pm l,0)$. The signed area $l\varepsilon$ of $T(l,\varepsilon)$ distinguishes mirror images $T(l,\pm\varepsilon)$ but is not Lipschitz continuous under perturbations. Indeed, as $\varepsilon\to 0$, the triangle $T(l,\varepsilon)$ degenerates to a straight line, while the area drops to 0 too quickly so that $\dfrac{\mathrm{vol}(T(l,\varepsilon))-\mathrm{vol}(T(l,0))}{\varepsilon-0}=\dfrac{l\varepsilon}{\varepsilon}=l$ is not bounded. Hence, if given points are not restricted to a fixed bounded region, a small change in their positions may lead to a large change for the area of a triangle, and similarly for the volume of a high-dimensional simplex in $\mathbb{R}^{n}$.
Problem 1.4 (Lipschitz continuous detection of degenerate simplices).
Find a Lipschitz continuous real-valued function $f(A)$ for all simplices $A$ on $n+1$ points $p_{0},\dots,p_{n}\in\mathbb{R}^{n}$ such that $f(A)=0$ if and only if $A$ is degenerate.
2 The strength is a Lipschitz continuous invariant
We solve Problem 1.4 by introducing the strength function whose Lipschitz continuity is proved in Theorem 2.8.
Definition 2.5 (the strength $\sigma(A)$ and signed strength $s(A)$ of a simplex $A\subset\mathbb{R}^{n}$).
Let $A\subset\mathbb{R}^{n}$ be the simplex on any $n+1$ points $p_{0},p_{1},\dots,p_{n}$ in $\mathbb{R}^{n}$. The half-perimeter $p(A)=\dfrac{1}{2}\sum\limits_{i\neq j}|p_{i}-p_{j}|$ is one half of the sum of all distances between the vertices of $A$. The strength of $A$ is defined as $\sigma(A)=\dfrac{\mathrm{vol}^{2}(A)}{p^{2n-1}(A)}$. The signed strength is $s(A)=\mathrm{sign}(A)\sigma(A)$.
Example 2.6 (the strength of a line segment).
For $n=1$, the simplex $A$ on any points $p_{0},p_{1}\in\mathbb{R}$ is the line segment with $\mathrm{vol}(A)=|p_{1}-p_{0}|=2p(A)$, strength $\sigma(A)=\dfrac{\mathrm{vol}^{2}(A)}{p(A)}=2|p_{1}-p_{0}|$, and signed strength $s(A)=2(p_{1}-p_{0})$.
Example 2.7 (the strength of a triangle).
(a) Let $A\subset\mathbb{R}^{2}$ be a triangle with sides $a,b,c$. Heron’s formula for the area $\mathrm{vol}(A)=\sqrt{p(p-a)(p-b)(p-c)}$, where $p=\dfrac{a+b+c}{2}=p(A)$ is the half-perimeter, gives the strength $\sigma(A)=\dfrac{\mathrm{vol}^{2}(A)}{p^{3}(A)}=\dfrac{(p-a)(p-b)(p-c)}{p^{2}}$. The triangle $T(l,\varepsilon)$ in Example 1.3 has $p=\dfrac{l+\varepsilon+\sqrt{l^{2}+\varepsilon^{2}}}{2}$. Using $l\leq\sqrt{l^{2}+\varepsilon^{2}}\leq l+\varepsilon$, the strength $\sigma=\dfrac{(l-\varepsilon+\sqrt{l^{2}+\varepsilon^{2}})(\varepsilon+\sqrt{l^{2}+\varepsilon^{2}}-l)(l+\varepsilon-\sqrt{l^{2}+\varepsilon^{2}})}{2(l+\varepsilon+\sqrt{l^{2}+\varepsilon^{2}})^{2}}\leq\dfrac{2l\cdot 2\varepsilon\cdot\varepsilon}{2(2l)^{2}}\leq\dfrac{\varepsilon^{2}}{2l}\leq\dfrac{\varepsilon}{2}$ for any $0\leq\varepsilon\leq l$. Hence, the strength of $T(l,\varepsilon)$ is Lipschitz continuous with respect to perturbations of $\varepsilon$.
(b) To visualize the strength $\sigma(A)$, we assume that $0<a\leq b\leq c$ and normalize $A$ by the larger side $c$ to get the sides $\tilde{a}=\dfrac{a}{c}\leq\tilde{b}=\dfrac{b}{c}\leq\tilde{c}=1$. The resulting space of normalized triangles is parametrized by the coordinates $x=\dfrac{a}{c}$ and $y=1-\dfrac{b}{c}$ in the triangular region $\Delta=\{(x,y)\in\mathbb{R}^{2}\mid x\in[0,1],\,x\geq y,\,x+y\leq 1\}$, where $x+y\leq 1$ means that $a\leq b$, while $x\geq y$ is equivalent to the triangle inequality $a+b\geq c$, see Fig. 1 (left). The half-perimeter is $\tilde{p}=\dfrac{1}{2}(\tilde{a}+\tilde{b}+\tilde{c})=\dfrac{1}{2}(x+(1-y)+1)=\dfrac{1}{2}(2+x-y)$. Then $\tilde{p}-\tilde{a}=\dfrac{1}{2}(2-x-y)$, $\tilde{p}-\tilde{b}=\dfrac{1}{2}(x+y)$, $\tilde{p}-\tilde{c}=\dfrac{1}{2}(x-y)$, and the strength is $\sigma(x,y)=\dfrac{(2-x-y)(x^{2}-y^{2})}{2(2+x-y)^{2}}$, see Fig. 1 (right). In the region $\Delta$, the horizontal side $\{x\in(0,1],\,y=0\}$ represents all (normalized) isosceles triangles with $\tilde{a}\leq\tilde{b}=\tilde{c}=1$ and strength $\sigma=\dfrac{(2-x)x^{2}}{2(2+x)^{2}}$ for $x\in[0,1]$. The right hand side $\{x\in(0.5,1],\,x+y=1\}$ of the region $\Delta$ represents all (normalized) isosceles triangles with $\tilde{a}=\tilde{b}\leq\tilde{c}=1$ and strength $\sigma=\dfrac{2x-1}{2(2x+1)^{2}}$ for $x\in[\frac{1}{2},1]$. The vertex $(x,y)=(1,0)$ represents all equilateral triangles with side 1 of strength $\sigma=\dfrac{1}{18}$. Any equilateral triangle with sides $a=b=c$ has the strength $\sigma=\dfrac{(a^{2}\sqrt{3}/4)^{2}}{(3a/2)^{3}}=\dfrac{a}{18}$.
Figure 1: Left: Example 2.7(b) parametrizes the space of (normalized) triangles with sides $0<a\leq b\leq c$ by $x=\dfrac{a}{c}$ and $y=1-\dfrac{b}{c}$ over the region $\Delta=\{(x,y)\in\mathbb{R}^{2}\mid x\in[0,1],\,x\geq y,\,x+y\leq 1\}$. Right: the strength $\sigma(x,y)=\dfrac{(2-x-y)(x^{2}-y^{2})}{2(2+x-y)^{2}}$ of a normalized triangle over the region $\Delta$.
Figure 2: Vertical sections of the strength $\sigma$ of normalized triangles parametrized by $x,y$ in Fig. 1. Left: isosceles triangles with $\tilde{a}=\tilde{b}\leq\tilde{c}=1$ have $y=0$ and $\sigma=\dfrac{(2-x)x^{2}}{2(2+x)^{2}}$ for $x\in[0,1]$. Right: isosceles triangles with sides $\tilde{a}\leq\tilde{b}=\tilde{c}$ have y=1xy=1-x and $\sigma=\dfrac{2x-1}{2(2x+1)^{2}}$ for $x\in[\frac{1}{2},1]$.
Recall that an isometry is any distance-preserving transformation of $\mathbb{R}^{n}$, which decomposes into translations and orthogonal maps from the group $\mathrm{O}(\mathbb{R}^{n})$. A rigid motion is any composition of translations and rotations from the special orthogonal group $\mathrm{SO}(\mathbb{R}^{n})$ The strength of a simplex was essentially used to define a Lipschitz continuous metric on invariants of $n$-dimensional clouds of $m$ unordered points, which are complete under rigid motion in $\mathbb{R}^{n}$ and can be computed in a polynomial time of $m$, for a fixed dimension $n$ [9, Theorem 4.7], see detailed proofs in [6].
Theorem 2.8 (strength properties: invariance, computational time, and Lipschitz continuity).
(a) The strength $\sigma(A)$ and signed strength $s(A)$ of any simplex $A\subset\mathbb{R}^{n}$ are invariant under isometry and rigid motion in $\mathbb{R}^{n}$, respectively, and can be computed in time $O(n^{3})$. The uniform scaling of $\mathbb{R}^{n}$ by any factor $c>0$ multiplies both $\sigma(A)$ and $s(A)$ by $c$.
(b) Fix any dimension $n\geq 1$. Then there is a constant $\lambda_{n}>0$ such that, for any $\varepsilon>0$, if a simplex $B$ is obtained from another simplex $A\subset\mathbb{R}^{n}$ by perturbing every vertex of $A$ within its $\varepsilon$-neighborhood, then $|\sigma(A)-\sigma(B)|\leq 2\lambda_{n}\varepsilon$, where $\lambda_{1}=2$, $\lambda_{2}=\sqrt{3}$, and $\lambda_{n}\leq\dfrac{2^{n+0.5}}{(n!)n^{2n-4}}$ for any $n\geq 3$, e.g. $\lambda_{3}<0.37$.
3 Proofs of strength properties in all dimensions
Proof of Theorem 2.8(b) for dimension $n=1$ and $\lambda_{1}=2$.
For $n=1$, a simplex $A\subset\mathbb{R}$ has two vertices $p_{0},p_{1}$ at a distance $d=|p_{0}-p_{1}|$ and the strength $\sigma(A)=2d$, which has Lipschitz constant $2\lambda_{1}=4$. Indeed, perturbing each of $p_{0},p_{1}$ up to $\varepsilon$ changes the distance $d$ up to $2\varepsilon$. ∎
We use vertical lines to denote the determinant $|M|$ of a matrix $M$ and the Euclidean length $|v|$ of a vector $v\in\mathbb{R}^{n}$.
Proof of Theorem 2.8(b) for dimension $n=2$ and $\lambda_{2}=\sqrt{3}$.
Let a triangle $A\subset\mathbb{R}^{2}$ have pairwise distances $a,b,c$. Using the half-perimeter $p=\dfrac{a+b+c}{2}$, the variables $\tilde{a}=p-a$, $\tilde{b}=p-b$, $\tilde{c}=p-c$ are expressed via $a,b,c$, so $a=\tilde{b}+\tilde{c}$, $b=\tilde{a}+\tilde{c}$, $c=\tilde{a}+\tilde{b}$, $p=\tilde{a}+\tilde{b}+\tilde{c}$. The Jacobian of this change of variables is $\left|\dfrac{\partial(a,b,c)}{\partial(\tilde{a},\tilde{b},\tilde{c})}\right|=\left|\begin{array}[]{ccc}0&1&1\\ 1&0&1\\ 1&1&0\end{array}\right|=2$.
If every point of $A$ is perturbed up to $\varepsilon$, then any pairwise distance between the vertices of $A$ changes by at most $2\varepsilon$. By the mean value theorem [5], this bound $2\varepsilon$ gives $|\sigma(A^{\prime})-\sigma(A)|\leq 2\varepsilon\sup|\nabla\sigma|$, where $\nabla\sigma=\left(\dfrac{\partial\sigma}{\partial a},\dfrac{\partial\sigma}{\partial b},\dfrac{\partial\sigma}{\partial c}\right)$ is the gradient of the first order partial derivatives of $\sigma(A)$ with respect to the three distances between points of $A$.
[…]
By the triangle inequalities for the sides $a,b,c$, we have $\max\{a,b,c\}\leq p=\dfrac{a+b+c}{2}$. Then $\dfrac{a}{p},\dfrac{b}{p},\dfrac{c}{p}\in(0,1]$, and $1-\dfrac{b}{p},1-\dfrac{c}{p}\in[0,1)$, and $2\dfrac{a}{p}-1\in(-1,1]$, so $\left|\dfrac{\partial\sigma}{\partial\tilde{a}}\right|\leq 1$. The similar bounds $\left|\dfrac{\partial\sigma}{\partial\tilde{b}}\right|,\left|\dfrac{\partial\sigma}{\partial\tilde{c}}\right|\leq 1$ imply that $|\nabla\sigma|\leq 2\sqrt{1^{2}+1^{2}+1^{2}}=2\sqrt{3}$, so $|\sigma(A^{\prime})-\sigma(A)|\leq 2\sup|\nabla\sigma|\varepsilon\leq 2\lambda_{2}\varepsilon$ for $\lambda_{2}=\sqrt{3}$. ∎
Theorem 2.8(b) for dimesions $n\geq 3$ will be proved by using Lemmas
Lemma 3.9 (edge ratios).
For any simplex $A$ on points $p_{0},\dots,p_{n}\in\mathbb{R}^{n}$, we have $\dfrac{|p_{i}-p_{j}|}{p(A)}\leq\dfrac{2}{n}$ for all $i,j=0,\dots,n$.
Proof.
Set $d_{ij}=|p_{i}-p_{j}|$ and use triangle inequalities: $2p(A)=\sum\limits_{k,l=0}^{n}d_{kl}\geq d_{ij}+\sum\limits_{k\neq i,j}(d_{ki}+d_{kj})\geq d_{ij}+(n-1)d_{ij}=nd_{ij}$. ∎
Definition 3.10 (Cayley-Menger determinant, [2, Chapter II, p.98]).
Let the simplex $A$ on any points $p_{0},\dots,p_{n}\in\mathbb{R}^{n}$ have the $(n+1)\times(n+1)$ matrix $D_{ij}$ of squared Euclidean distances $d_{ij}^{2}=|p_{i}-p_{j}|^{2}$ for $i,j=0,\dots,n$. The $(n+2)\times(n+2)$ matrix $\hat{D}$ is obtained from $D$ by adding the top row $(0,1,\dots,1)$ and the left column $(0,1,\dots,1)^{T}$. The Cayley-Menger determinant [3, 7] expresses the squared volume of the simplex AA as $\mathrm{vol}^{2}(A)=\dfrac{(-1)^{n-1}}{2^{n}(n!)^{2}}\det\hat{D}$.
Proof of Theorem 2.8(a).
Any isometry preserves all distances and hence the strength $\sigma(A)$ expressed via distances in Definition 2.5. Any special orthogonal matrix $M\in\mathrm{SO}(\mathbb{R}^{n})$ keeps the sign of a simplex $A$, i.e. multiplies $\mathrm{sign}(A)$ by $\det M=1$. Then any rigid motion preserves the signed strength $s(A)=\mathrm{sign}(A)\sigma(A)$. To compute $\sigma(A)$ and $s(A)$, we need the half-perimeter $p(A)$, which requires a quadratic time $O(n^{2})$, and then $\mathrm{sign}(A)$ and $\mathrm{vol}^{2}(A)$ by using determinants of sizes up to $n+2$, which can be calculated in time $O(n^{3})$ by matrix diagonalization [8, section 11.5]. ∎
The rencontre number $r_{n}=n!\sum\limits_{k=0}^{n}\dfrac{(-1)^{k}}{k!}$ counts all permutations of $1,\dots,n$ without a fixed point [4] and equals the integer nearest to $\dfrac{n!}{e}$, e.g. $r_{2}=1$, $r_{3}=2$, $r_{4}=9$, $r_{5}=44$.
Lemma 3.11.
In the notations of Definition 3.10, we have $\dfrac{\det\hat{D}}{p^{2n}(A)}\leq r_{n+2}\left(\dfrac{2}{n}\right)^{2n}$ for any $n\geq 1$.
Proof.
The determinant formula $\det\hat{D}=\sum\limits_{\xi\in S_{n+2}}(-1)^{\mathrm{sign}(\xi)}\hat{D}_{1,\xi(1)}\dots\hat{D}_{n,\xi(n)}$ over all permutations $\xi\in S_{n+2}$ excludes all zeros on the diagonal. Then $k\neq\xi(k)$ for $k=0,\dots,n+1$, i.e. we can consider only permutations $\xi$ of $0,\dots,n+1$ that have no fixed elements. The number of such permutations is the rencontre number $r_{n+2}$. Then $\det\hat{D}$ is a sum of $r_{n+2}$ non-zero terms, each being a product of $n$ squared distances $d_{ij}$. The upper bound $d_{ij}\leq\dfrac{2}{n}p(A)$ in Lemma 3.9 implies that each term in the sum of $\det\hat{D}$ is at most $\left(\dfrac{2}{n}p(A)\right)^{2n}$. Then $\dfrac{\det\hat{D}}{p^{2n}(A)}\leq r_{n+2}\left(\dfrac{2}{n}\right)^{2n}$ as required. ∎
For $n=1$ in Lemma 3.11, we have $\det\hat{D}=2d_{01}^{2}$ and $p(A)=\dfrac{d_{01}}{2}$, so $\dfrac{\det\hat{D}}{p^{2}(A)}=8\leq r_{3}(2^{2})=8$ as expected.
Lemma 3.12.
In the notations of Definition 3.10, we have $\left|\dfrac{\partial\det\hat{D}}{\partial d_{ij}}\right|\dfrac{1}{p^{2n-1}(A)}\leq 4(r_{n}+r_{n+1})\left(\dfrac{2}{n}\right)^{2n-1}$ for any fixed $i\neq j$.
Proof.
Since $\det\hat{D}$ has $d_{ij}^{2}$ in exactly two cells in different rows and columns, $\det\hat{D}$ is a quadratic polynomial $\alpha d_{ij}^{4}+\beta d_{ij}^{2}+\gamma$ for some $\alpha,\beta,\gamma$ that depend on other fixed distances $d_{kl}\neq d_{ij}$. Then $\dfrac{\partial\det\hat{D}}{\partial d_{ij}}=4\alpha d_{ij}^{3}+2\beta d_{ij}$.
The coefficient $\alpha$ is the determinant of the $n\times n$ submatrix $(b_{ij})$ obtained from $\hat{D}$ by removing two rows and columns indexed by $i+2,j+2$. For example, fix $i=0$ and $j=1$. If $n=2$, then $\alpha=\left|\begin{array}[]{cc}0&1\\ 1&0\end{array}\right|=-1$. If $n=3$, then $\alpha=\left|\begin{array}[]{ccc}0&1&1\\ 1&0&d_{23}^{2}\\ 1&d_{32}^{2}&0\end{array}\right|=2d_{23}^{2}$. Since the matrix $(b_{ij})$ has zeros on the main diagonal, its determinant $\alpha=\sum\limits_{\xi\in S_{n}}(-1)^{\mathrm{sign}(\xi)}b_{1,\xi(1)}\dots b_{n,\xi(n)}$ is a sum over all permutations $\xi\in S_{n}$ with no fixed points. Then the sum $\alpha$ has $r_{n}$ non-zero products $b_{1,\xi(1)}\dots b_{n,\xi(n)}$ and the total degree 2(n2)2(n-2) in all distances $d_{kl}\neq d_{ij}$. After dividing the polynomial $\alpha d_{ij}^{3}$ of the degree 2n12n-1 by $p^{2n-1}(A)$, we use the upper bound $\dfrac{d_{kl}}{p(A)}\leq\dfrac{2}{n}$ in Lemma 3.9 to get $\dfrac{|\alpha|d_{ij}^{3}}{p^{2n-1}(A)}\leq r_{n}\left(\dfrac{2}{n}\right)^{2n-1}$.
The coefficient $\beta$ in $\det\hat{D}=\alpha d_{ij}^{4}+\beta d_{ij}^{2}+\gamma$ is the sum of products from the determinants of the two submatrices obtained from $\hat{D}$ by removing row $i+2$ and column $j+2$ (for one submatrix), then row $j+2$ and column $i+2$ (for another submatrix). If $n=3$, $i=0$, $j=2$, the determinants are $\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{10}^{2}&0&d_{13}^{2}\\ 1&d_{20}^{2}&d_{21}^{2}&d_{23}^{2}\\ 1&d_{30}^{2}&d_{32}^{2}&0\end{array}\right|+\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{01}^{2}&d_{02}^{2}&d_{03}^{2}\\ 1&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$.
Since each $(n+1)\times(n+1)$ submatrix includes one entry $d_{ij}^{2}$, we exclude all products with this entry, which are multiplied by the removed $d_{ij}^{2}$ from $\hat{D}$ and hence were counted in $\alpha d_{ij}^{4}$. Hence we can replace $d_{02}=d_{02}$ with 0 and get $\beta=\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{10}^{2}&0&d_{13}^{2}\\ 1&0&d_{21}^{2}&d_{23}^{2}\\ 1&d_{30}^{2}&d_{32}^{2}&0\end{array}\right|+\left|\begin{array}[]{cccc}0&1&1&1\\ 1&d_{01}^{2}&0&d_{03}^{2}\\ 1&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$. Up to a permutation of indices, each submatrix can be rewritten with the diagonal that has one original $d_{ij}^{2}$ (now replaced with 0), while all other diagonal elements are initial zeros. The example above gives $\beta=-\left|\begin{array}[]{cccc}0&1&1&1\\ 1&0&d_{21}^{2}&d_{23}^{2}\\ 1&d_{10}^{2}&0&d_{13}^{2}\\ 1&d_{30}^{2}&d_{32}^{2}&0\end{array}\right|-\left|\begin{array}[]{cccc}0&1&1&1\\ 1&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{01}^{2}&0&d_{03}^{2}\\ 1&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$. Similarly to the argument for the determinant $\alpha$, the sum $\beta d_{ij}$ contains $2r_{n+1}$ products of the total degree 2n12n-1, so $\dfrac{|\beta|d_{ij}}{p^{2n-1}(A)}\leq 2r_{n+1}\left(\dfrac{2}{n}\right)^{2n-1}$ by Lemma 3.9. Then the required inequality follows: $\left|\dfrac{\partial\det\hat{D}}{\partial d_{ij}}\right|\dfrac{1}{p^{2n-1}(A)}=\dfrac{|4\alpha d_{ij}^{3}+2\beta d_{ij}|}{p^{2n-1}(A)}\leq(4r_{n}+4r_{n+1})\left(\dfrac{2}{n}\right)^{2n-1}$. ∎
Proof of Theorem 2.8(b) for $n\geq 3$.
For $n=3$, the squared volume is $\mathrm{vol}^{2}(A)=\dfrac{1}{288}\left|\begin{array}[]{ccccc}0&1&1&1&1\\ 1&0&d_{01}^{2}&d_{02}^{2}&d_{03}^{2}\\ 1&d_{10}^{2}&0&d_{12}^{2}&d_{13}^{2}\\ 1&d_{20}^{2}&d_{21}^{2}&0&d_{23}^{2}\\ 1&d_{30}^{2}&d_{31}^{2}&d_{32}^{2}&0\end{array}\right|$.
Similarly to the case $n=2$, the mean value theorem [5] for the strength $\sigma(A)=\dfrac{\mathrm{vol}^{2}(A)}{p^{2n-1}(A)}$ implies that $|\sigma(A)-\sigma(B)|\leq 2\varepsilon\sup|\nabla\sigma|\leq 2\varepsilon\sqrt{\sum\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|^{2}}\leq 2\varepsilon\sqrt{\dfrac{n(n+1)}{2}}\max\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|.$ To find an upper bound of $\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|$, we initially ignore the numerical factor in the square volume $\mathrm{vol}^{2}(A)=\dfrac{(-1)^{n-1}}{2^{n}(n!)^{2}}\det\hat{D}$ and differentiate $\det\hat{D}\cdot\dfrac{1}{p^{2n-1}(A)}$ by the product rule: $\dfrac{\partial}{\partial d_{ij}}\left(\dfrac{\det\hat{D}}{p^{2n-1}(A)}\right)=\dfrac{\partial\det\hat{D}}{\partial d_{ij}}\cdot\dfrac{1}{p^{2n-1}(A)}-\dfrac{\det\hat{D}}{p^{2n}(A)}\cdot\dfrac{2n-1}{2}$. Lemmas 3.11 and 3.12 below imply the upper bound $\left|\dfrac{\partial}{\partial d_{ij}}\left(\dfrac{\det\hat{D}}{p^{2n-1}(A)}\right)\right|\leq(4r_{n}+4r_{n+1})\left(\dfrac{2}{n}\right)^{2n-1}+\left(n-\dfrac{1}{2}\right)r_{n+2}\left(\dfrac{2}{n}\right)^{2n}<\left(4r_{n}+4r_{n+1}+2r_{n+2}\right)\left(\dfrac{2}{n}\right)^{2n-1}$. Taking into account the factors $\dfrac{(-1)^{n-1}}{2^{n}(n!)^{2}}$ in $\mathrm{vol}^{2}(A)$ and $\sqrt{\dfrac{n(n+1)}{2}}$ for estimating the length of the gradient $\nabla\sigma$ of $\dfrac{n(n+1)}{2}$ first order partial derivatives $\dfrac{\partial\sigma}{\partial d_{ij}}$, we get $|\sigma(A)-\sigma(B)|\leq 2\varepsilon\sqrt{\dfrac{n(n+1)}{2}}\max\limits_{i\neq j}\sup\left|\dfrac{\partial\sigma}{\partial d_{ij}}\right|\leq 2\varepsilon c_{n}$ for the upper bound $c_{n}=\dfrac{\sqrt{n(n+1)/2}}{2^{n}(n!)^{2}}\left(4r_{n}+4r_{n+1}+2r_{n+2}\right)\left(\dfrac{2}{n}\right)^{2n-1}=(2r_{n}+2r_{n+1}+r_{n+2})\dfrac{2^{n-0.5}\sqrt{n+1}}{(n!)^{2}n^{2n-1.5}}$. The estimate $r_{n}\leq\dfrac{n!}{2}$ gives $c_{n}<\dfrac{n!}{2}(2+2(n+1)+(n+1)(n+2))\dfrac{2^{n-0.5}\sqrt{n+1}}{(n!)^{2}n^{2n-1.5}}=(n^{2}+5n+6)\sqrt{n+1}\dfrac{2^{n-1.5}}{(n!)n^{2n-1.5}}=\left(1+\dfrac{5}{n}+\dfrac{6}{n^{2}}\right)\sqrt{1+\dfrac{1}{n}}\dfrac{2^{n-1.5}}{(n!)n^{2n-4}}$. For $n\geq 3$, we get $1+\dfrac{5}{n}+\dfrac{6}{n^{2}}\leq\dfrac{10}{3}$ and $\sqrt{1+\dfrac{1}{n}}\leq\dfrac{2}{\sqrt{3}}$, so $c_{n}<\dfrac{20}{3\sqrt{3}}\cdot\dfrac{2^{n-1.5}}{(n!)n^{2n-4}}=\dfrac{5\sqrt{2}}{3\sqrt{3}}\dfrac{2^{n}}{(n!)n^{2n-4}}<\dfrac{2^{n+0.5}}{(n!)n^{2n-4}}=b_{n}$. ∎